OpenStudy (anonymous):

Suggest one reason why the fringe pattern produced by light passing through a diffraction grating is brighter than that produced from the same source with a double slit

OpenStudy (samgrace):

The angular width of the interference maximum, which is equal to 2 * theta_mins, is inversely proportional to N. Therefore the greater the number of illuminate slits N, the sharper the maximum. Diffracton grating: \[Nd \sin \theta = \lambda \] where N is the number of slits

OpenStudy (anonymous):

Isn't N the order of the diffracted beam? @samgrace

OpenStudy (samgrace):

No

OpenStudy (samgrace):

dsin(theta) = m*lamda , where m = 1, 2 ,3 , m is the order of diffration

OpenStudy (anonymous):

Okay, N is the number of slits, right. So the distance between slits and number of slits is related by d=1/N

OpenStudy (anonymous):

If N is greater, d decreases.

OpenStudy (samgrace):

Yeah

OpenStudy (anonymous):

And thus if d decreases, the order (m) also decreases. So, the lights is less spread out and brighter?

OpenStudy (samgrace):

The first and the last slits are separated by approximately N*d, so the intensity will be zero at angle theta_min, I think correct.